Is the graph of y=sin(x5) increasing or decreasing when x=13?
(enter increasing, decreasing, or neither).

Is it concave up or concave down?
1
Expert's answer
2020-03-03T11:30:23-0500

let's find a first derivative


"y'=(sin(x^5))'=(sin(x^5))'*(x^5)'=5*x^4*cos(x^5)"


now we have to understand whether the graph y' increasing or decreasing. So let's find the value of the derivative y' at x=13


"y'=5*x^4*cos(x^5)=5*13^4*cos(13^5)=106353\implies y'>0" ,


so the graph is increasing at x=13.


Next question - is the graph concave up or concave down. Let's find second derivative y"


"y''=(sin(x^5))''=(5*x^4*cos(x^5))'=(x^4)'*cos(x^5)+x^4*(cos(x^5))'=4*x^3*cos(x^5)+x^4*(-5*x^4*sin(x^5))=-25*x^8*sin(x^5)+20*x^3*cos(x^5)"


to understand is the graph concave up or concave down let's find y'' at x=13


"y''=-25*x^8*sin(x^5)+20*x^3*cos(x^5)=5*x^3*(4*cos(x^5)-5*x^3*sin(x^5))"

"y''= -13609384564.031\implies y''<0"


so the graph is concave down